OpenStudy (anonymous):

Jeannie works 5 hours on Tuesday, 3 hours on Wednesday, and 4.5 hours on Thursday. If h represents her hourly pay rate, which expression below shows the amount she earned?

OpenStudy (anonymous):

Can I see the expressions?

OpenStudy (anonymous):

12.5 12.5h 12.5+h 12.5/5

OpenStudy (anonymous):

B

OpenStudy (anonymous):

The formula for the area of a triangle is A=1/2bh. Which is the equivalent equation solved for h?

OpenStudy (anonymous):

It's saying that Jen makes $12.5 times a number that represents the number of hours she works. You're complicating things going at it the way you are.

OpenStudy (anonymous):

2A/b=h 1/2ab=h 2ab=h A/2B=h

OpenStudy (anonymous):

Thats whole new Question

OpenStudy (anonymous):

Ummm lol wrong day to ask me, just got a concussion so math requiring me to think harder is more complicated than normal lol

OpenStudy (anonymous):

What is the simplified expression for the expression below? 7(x-4)-3(x+5)

OpenStudy (anonymous):

7 - 3x